OpenStudy (anonymous):

Complete the square to solve the equation below. x^2+x=19/4

OpenStudy (anonymous):

hey buddy can you help please

OpenStudy (anonymous):

\[adding both sides \left( \frac{ co-efficient of x }{2 } \right)^{2}=\left( \frac{ 1 }{2} \right)^{2} =\frac{ 1 }{4 }\]

OpenStudy (anonymous):

1/4 isnt the answer

OpenStudy (anonymous):

i have given you the hint.

OpenStudy (anonymous):

can you proceed further

OpenStudy (anonymous):

\[x=-\frac{ 1 }{ 2}+\sqrt{5}\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[x=-\frac{ 1 }{ 2 }\pm \sqrt{5}\]

OpenStudy (anonymous):

oh okay i was right except for the +/-

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw